# On the Consistency of GNN Explanations for Malware Detection

### v1

@@ Abstract
Index Terms:
Graph Neural Network, Explainability, Machine Learning, Malware Detection, Dynamic Analysis, Control Flow Graph.I Introduction
In the past few years, malware attacks have escalated dramatically, highlighting the limitations of conventional malware detection methods, such as signature-based techniques. While these methods are quick and widely utilized, they fall short in identifying zero-day and sophisticated malware threats [1]. Consequently, there has been a significant shift towards integrating machine learning (ML) strategies, noted for their enhanced detection capabilities, adaptability to emerging threats, and reduced false positives. Numerous studies have thus shifted focus towards leveraging ML [2, 3, 4, 5] for malware detection and classification.
As malware becomes more complex and developers employ advanced tactics to evade detection, the need for robust representation of malware samples has become evident. Recent research highlights the effectiveness of graph-based models in depicting malware behavior, which facilitates improved detection performance. These models provide a detailed view of a program’s execution, helping analysts understand the program’s logic, pinpoint vulnerabilities, and expose malicious activities, including hidden or scrambled code. Various graph structures, including Control Flow Graph (CFG) [6, 7, 8, 9], Function Call Graph (FCG) [10, 11, 12], and Application Programming Interface (API) Call Graph (ACG) [13, 14, 15], have been increasingly utilized to feed data into ML models designed for malware detection. Although initial efforts involved using these graph structures with CNN or RNN architectures, the advent of Graph Neural Networks (GNNs) with variants like Graph Convolutional Networks (GCNs), Graph Isomorphism Networks (GINs), GraphSage, and Graph Attention Networks (GATs) have demonstrated superior outcomes when integrated with CFGs and FCGs.
@@
The structure of the paper is organized as follows: Section II provides a review of the relevant background. Section III details the proposed framework, while Section IV focuses on the explainability metrics in more detail. Section V, presents the experimental results and analysis, Finally, Section VI concludes the paper and outlines potential directions for future research.
II Background
@@
An alternative static detection model for malicious JavaScript is introduced in JStrong [18]. It generates an abstract syntax tree from JavaScript source code and integrates data flow and control flow information into a program dependency graph. A GNN is then employed to analyze the graph and classify malicious code based on its structural patterns.
@@
III Proposed Method
@@
III-A Node Feature Embedding
The dynamically generated CFGs contain several types of features in each node, but not all are suitable or useful for the decision-making module. The assembly instructions of each node in the CFG are selected as the key features, and a two-step embedding technique is used to map them into a 64-dimensional real-valued vector. This process includes a rule-based instruction encoding strategy followed by machine learning-based dimensionality reduction using an autoencoder, as shown in Figure 2.
III-A1 Rule-Based Instruction Encoding
To encode the assembly instructions of CFG nodes, we adapted the encoding process outlined in [7] with slight modifications to better suit the structure of dynamically generated CFGs. Each x86-64 assembly instruction consists of up to seven components: option, prefix, opcode, ModRM, SIB, displacement, and immediate. The encoding process for each component is defined as follows:
Prefix: The prefix includes four fields: the extra segment (ES) register, operand-size override, address-size override, and lock prefix. The ES segment register has seven possible values, while the other three fields are binary (0 or 1). Thus, the prefix is encoded as a nine-dimensional one-hot vector.
@@ III-A1 Rule-Based Instruction Encoding
The final encoded vector for each instruction is formed by concatenating all these components, resulting in a 438-dimensional vector.
@@
The high-dimensional 438-dimensional vectors are reduced to a 64-dimensional latent representation using an autoencoder. The autoencoder consists of an encoder-decoder architecture, where the encoder reduces the dimensionality, and the decoder reconstructs the original vector. The reduced representation retains the most relevant information for malware detection.
@@
where is the number of training samples, is the encoder function, and is the decoder function. Once trained, the encoder reduces the 438-dimensional instruction vector into a compact 64-dimensional feature vector used as the final node feature embedding for the GNN model:
@@
III-B Graph Classification Using GNNs
@@
III-B1 Node Embedding with GNN Layers
@@
where is the embedding of node at layer , represents the initial node feature vector, is the set of neighboring nodes of node , is an aggregation function that collects information from neighboring nodes, and is an update function that combines the aggregated information with the node’s current embedding.
The aggregation function and update function can vary depending on the specific GNN model. For instance, in the case of Graph Convolutional Networks (GCNs), the aggregation is based on a normalized summation of neighboring node features, and the update step applies a linear transformation followed by a non-linearity:
@@ III-B2 Readout Layer
III-B3 Graph Classification
@@ III-C GNN Explanation Techniques
where is the class label, is the predicted label, is the adjacency matrix of the input graph, is the learnable mask that assigns a weight to each edge of the input graph based on its importance for the model’s prediction, is the sigmoid function that maps the weight to the range , and is the corresponding node feature set. This optimization is performed separately for each input sample to generate its corresponding explanation. After optimization, the top-ranked edges based on their weights are selected to form the important subgraph.
PGExplainer uses the following cross-entropy loss to train the explainer on multiple samples. Once trained, the explainer model can be applied to other samples without requiring retraining. For binary classification, the objective is formulated as:
where is the original input graph for sample , is the sampled explanatory subgraph for sample in the -th sampling step, and represents the parameters of the explainer model.
@@
III-D RankFusion Explainer: An Edge Ranking Aggregation Strategy
To enhance the accuracy of explainers, we propose a novel aggregation method called RankFusion explainer that combines the edge rankings from two explainers. The goal is to improve the quality of explanations across different sparsity levels.
The RankFusion process involves reassigning edge weights to their descending rank order for each explainer independently. For each pair of similar edges present in both explainers, we compute the absolute difference between their ranks. If this difference is smaller than a predefined threshold, set as a percentage of the total number of edges in the graph, we assign the rank of the edge as the maximum of the two ranks. However, if the difference exceeds the threshold, the rank of the edge is assigned based on the rank from the explainer with the highest accuracy at the corresponding explanation percentage.
To evaluate the RankFusion explainer, we retain a certain percentage of edges with the highest aggregated weights, reconstruct the graph, and feed it into a trained GNN model. The accuracy of the GNN on this reduced graph serves as the performance measure of the RankFusion explainer. This strategy ensures that the RankFusion explainer retains the most meaningful edges, combining the strengths of both explainers to produce more consistent and accurate explanations across varying sparsity levels. The RankFusion algorithm is presented in Algorithm 1.
IV Explainer Subgraph Evaluation and Construction
@@
where is the important subgraph, is the original input, and is the trained GNN model. measures the prediction difference between the original input and its unimportant part, capturing the contribution of the removed edges to the model’s decision. Conversely, compares the prediction between the original graph and the important subgraph, reflecting the predictive capacity of the extracted subgraph.
@@
Each perturbed graph is retained for consistency evaluation if it meets the following conditions:
The model prediction is unchanged:
(13) -
The graph-level embeddings remain similar:
(14) where denotes the model’s embedding function, and is a threshold for cosine similarity.
@@ IV-1 Top-Edge Selection (TES)
IV-2 Greedy Edge-wise Composition (GEC)
@@
By prioritizing both edge weight and structural connectivity, GEC ensures that the selected subgraph retains the most informative and influential components of the original graph. This results in more faithful and robust explanations, ultimately supporting a more reliable interpretation of the model’s decision-making process.
V Results and Analysis
For our experiments, we selected 1,117 malicious samples from BODMAS [30], 1,029 malicious samples from PMML [31], and 510 benign samples from DikeDataset [32], providing a balanced dataset for evaluating the effectiveness of the explainers across various malware families and sample types. To recover CFGs dynamically, we employ the angr library [33, 34, 35], a Python-based binary analysis tool. angr constructs graphs by integrating both symbolic execution and constraint solving, enabling comprehensive and precise CFG recovery. The dynamically generated CFGs used in our experiments are available for public access here.
We use a three-layer GCN with a 20% dropout rate for malware detection. The detection model achieved strong performance, with an accuracy of 94.74% and an F1 score of 96.82%. These results highlight the effectiveness of the proposed framework, which combines dynamically generated CFGs, node feature embeddings based on assembly instructions, and a GNN-based detection model to accurately differentiate between benign and malicious samples.
@@ V Results and Analysis
We assess the performance of the explainers using the Fidelity-, Fidelity+, and Consistency metrics across different sparsity levels. Figure 4 presents the Fidelity- and Fidelity+ results under the GEC approach. In each subplot, the x-axis represents the sparsity level, defined as the percentage of edges retained from the original graph, and the y-axis indicates the fidelity score. The solid lines show the fidelity values computed from the original, unperturbed graphs for each explainer and sparsity level.
@@
To assess consistency in practice, we generated 20 perturbed versions of each input graph: 10 with random node removal and 10 with random edge removal. In our implementation, the perturbation ratio was set to . Each perturbed graph was retained for analysis only if it preserved the model’s original prediction and the cosine similarity between the graph-level embeddings of the original and perturbed graphs was below a selected threshold. We performed a sweep over similarity thresholds ranging from 0.01 to 0.03 in increments of 0.0025 to identify valid perturbations. For each explainer, we visualized the variability of Fidelity+ and Fidelity- scores across all valid perturbations. The solid lines represent fidelity values computed on the original (unperturbed) graphs, while the shaded regions reflect the range between the minimum and maximum fidelity scores observed across the valid perturbations.
@@
The results show that Fidelity- values are generally low and stable for explainers, when subgraphs are extracted using GEC. This is particularly evident for GNNExplainer, whose performance stabilizes significantly under this approach. According to the definition of Fidelity-, lower values are desirable, and this trend is clearly observed for the best-performing explainers (IG, GBP, and SAL), which consistently yield lower Fidelity- scores.
@@
In contrast, Fidelity+ values demonstrate a consistent pattern across sparsity levels. The top explainers (IG, GBP, and SAL) achieve higher scores, indicating a substantial change in model prediction when the important subgraph is removed. This suggests that the identified important subgraph indeed contains the critical predictive structure. Furthermore, the consistency bands in the Fidelity+ plot are generally narrow, indicating that explanations remain stable across perturbations. Although Fidelity- bands are somewhat wider, the results still indicate reasonable stability, particularly for the top explainers. These findings highlight the advantage of using GEC in conjunction with effective explainers for generating reliable and robust explanations.
@@
To assess the effectiveness of the proposed RankFusion explainer, we investigate whether selecting and combining the outputs of the top-performing explainers can lead to improved subgraph quality and model performance. Based on the previous analysis across accuracy, fidelity, and consistency metrics, the two most effective explainers identified were Integrated Gradients and Guided Backpropagation. RankFusion is applied by aggregating the edge importance rankings produced by these two explainers using the methodology described earlier.
@@
Figure 5 compares the model’s accuracy when using subgraphs generated by IG, GBP, and their aggregated result. All subgraphs in this evaluation are extracted using the GEC method. The x-axis represents the sparsity level, and the y-axis shows the accuracy obtained by feeding the resulting subgraph into the pre-trained GCN. Across all sparsity levels, the aggregated explanation consistently achieves higher accuracy than either of the individual explainers. While the improvement over the best individual explainer (IG) is modest, the results clearly support the hypothesis that selecting and combining the outputs of strong explainers leads to more informative subgraphs.
@@
VI Conclusion
This paper presents a comprehensive framework for malware detection using dynamically generated CFGs. Node-level assembly instructions are embedded into a continuous vector space through an embedding approach that combines rule-based encoding with autoencoder-based learning. These representations are then used as input to a GNN-based classifier for malware detection. To enhance the interpretability of the model, five explainers are employed: GNNExplainer, PGExplainer, and CaptumExplainer with three attribution methods, namely Integrated Gradients, Guided Backpropagation, and Saliency. In addition, we propose RankFusion, an aggregated explainer that selects and integrates the outputs of the top-performing explainers to improve the quality of generated explanations. We also introduce GEC as a connectivity-aware subgraph extraction strategy that mitigates the shortcomings of traditional top-edge selection. Experimental results demonstrate the strong detection performance of the framework, with an accuracy exceeding 94% and an F1 score above 96%. Further analysis based on Fidelity, Accuracy, and Consistency metrics shows that the combination of RankFusion and GEC produces more stable, interpretable, and accurate explanations, reinforcing the value of incorporating both explainer aggregation and structurally coherent subgraph extraction in GNN-based malware detection.

### v2

@@ Abstract
keywords:
[…]
In the past few years, malware attacks have escalated dramatically, highlighting the limitations of conventional malware detection methods, such as signature-based techniques. While these methods are quick and widely utilized, they fall short in identifying zero-day and sophisticated malware threats [1]. Consequently, there has been a significant shift towards integrating machine learning (ML) strategies, noted for their enhanced detection capabilities, adaptability to emerging threats, and reduced false positives. Numerous studies have thus shifted focus towards leveraging ML [2, 3] for malware detection and classification.
As malware becomes more complex and developers employ advanced tactics to evade detection, the need for robust representation of malware samples has become evident. Recent research highlights the effectiveness of graph-based models in depicting malware behavior, which facilitates improved detection performance. These models provide a detailed view of a program’s execution, helping analysts understand the program’s logic, pinpoint vulnerabilities, and expose malicious activities, including hidden or scrambled code. Various graph structures, including Control Flow Graph (CFG) [4, 5], Function Call Graph (FCG) [6, 7], and Application Programming Interface (API) Call Graph (ACG) [8, 9], have been increasingly utilized to feed data into ML models designed for malware detection. Although initial efforts involved using these graph structures with CNN or RNN architectures, the advent of Graph Neural Networks (GNNs) with variants like Graph Convolutional Networks (GCNs) [10], Graph Isomorphism Networks (GINs) [11], GraphSAGE [12], and Graph Attention Networks (GATs) [13] have demonstrated superior outcomes when integrated with CFGs and FCGs.
@@
The structure of the paper is organized as follows: Section 2 provides a review of the relevant background. Section 3 details the proposed framework, while Section 4 focuses on the explainability metrics in more detail. Section 5, presents the experimental results and analysis. Finally, Section 6 concludes the paper and outlines potential directions for future research.
2 Background
@@
Recent efforts have also explored advanced reinforcement learning strategies to improve intrusion detection in complex environments such as the Industrial Internet of Things (IIoT). Notable among these are heuristic and open-set intrusion detection systems leveraging Deep Q-Networks (DQN) and its variants for handling zero-day attacks and unknown threats with limited labeled data [25]. Additionally, real-time defense models such as RT-A3C enhance responsiveness in edge-enabled IIoT scenarios [26], while positive sampling techniques in self-supervised graph contrastive learning offer promising directions for robust graph-based representations [27].
3 Proposed Method
@@
3.1 Node Feature Embedding
The dynamically generated CFGs contain several types of features in each node (basic block), but not all are suitable or useful for the decision-making module. The assembly instructions of each node in the CFG are selected as the key features, and a two-step embedding technique is used to map them into a 64-dimensional real-valued vector. This process includes a rule-based instruction encoding strategy followed by machine learning-based dimensionality reduction using an autoencoder, as shown in Figure 2.
3.1.1 Rule-Based Instruction Encoding
To encode the assembly instructions of CFG nodes, we adapted the encoding process outlined in [4] with slight modifications to better suit the structure of dynamically generated CFGs. Our approach applies a fine-grained rule-based encoding scheme that provides a unique and full-coverage mapping for the x86-64 instruction set, ensuring robust handling of all possible instructions. This design inherently addresses the out-of-vocabulary problem by mapping any instruction to a fixed 439-dimensional vector. Each x86-64 assembly instruction consists of up to seven components: option, prefix, opcode, ModRM, SIB, displacement, and immediate. The encoding process for each component is defined as follows:
Prefix: The prefix includes four fields: the extra segment (ES) register, operand-size override, address-size override, and lock prefix. The ES segment register has seven possible values, while the other three fields are binary (0 or 1). Thus, the prefix is encoded as a ten-dimensional one-hot vector.
@@ 3.1.1 Rule-Based Instruction Encoding
The final encoded vector for each instruction is formed by concatenating all these components, resulting in a 439-dimensional vector.
@@
The high-dimensional 439-dimensional vectors are reduced to a 64-dimensional latent representation using an autoencoder. The autoencoder consists of an encoder-decoder architecture, where the encoder reduces the dimensionality, and the decoder reconstructs the original vector. The reduced representation retains the most relevant information for malware detection.
@@
where is the number of training samples, is the encoder function, and is the decoder function. Once trained, the encoder reduces the 439-dimensional instruction vector into a compact 64-dimensional feature vector used as the final node feature embedding for the GNN model:
@@
This architecture offers key advantages in our context. Since only the entire CFG is labeled and individual basic blocks (nodes) do not have labels, end-to-end training approaches such as a Multi-Layer Perceptron (MLP) are not suitable. The autoencoder enables unsupervised dimensionality reduction, allowing the model to learn meaningful node representations without requiring node-level supervision. Additionally, attempting to jointly train a GNN with raw 439-dimensional input features and the graph structure may introduce instability during optimization and increase the risk of overfitting, particularly in security-sensitive domains. The autoencoder thus serves as an effective intermediate step that balances expressive representation with training stability.
3.2 Graph Classification Using GNNs
@@
3.2.1 Node Embedding with GNN Layers
@@
where is the embedding of node at layer , represents the initial node feature vector, is the set of neighboring nodes of node , and is an update function that combines the aggregated information with the node’s current embedding.
The aggregation function and update function can vary depending on the specific GNN model. For instance, in the case of GCNs, the aggregation is based on a normalized summation of neighboring node features, and the update step applies a linear transformation followed by a non-linearity:
@@ 3.2.2 Readout Layer
The rule-based instruction encoding is used to ensure robustness, full instruction coverage, and avoidance of out-of-vocabulary issues by mapping any valid instruction to a fixed-length 439-dimensional vector. This vector representation is semantically expressive, preserving the key structure and meaning of x86-64 instructions. The aggregation step enables flexible handling of varying numbers of instructions in each node. The dimensionality reduction through a trained autoencoder compresses the high-dimensional features into compact 64-dimensional embeddings, while preserving the most relevant information for the downstream classification task. This step significantly reduces computational overhead for GNN models without compromising the expressiveness of node features. The overall node feature embedding process, from rule-based instruction encoding to dimensionality reduction via autoencoder, is summarized in Algorithm 1.
3.2.3 Graph Classification
@@ 3.3 GNN Explanation Techniques
where is the class label, is the predicted label, is the adjacency matrix of the input graph, is the learnable mask that assigns a weight to each edge of the input graph based on its importance for the model’s prediction, is the sigmoid function that maps the weight to the range , and is the corresponding node feature set. The indicator function equals if , and otherwise. Moreover, denotes the prediction probability distribution output by the trained GNN model for a given input graph and node features . This optimization is performed separately for each input sample to generate its corresponding explanation. GNNExplainer jointly learns an edge mask and a node feature mask to identify the most influential structural and attribute components of the graph. After optimization, the top-ranked edges based on their weights are selected to form the important subgraph. The overall process of GNNExplainer is illustrated in Figure 3.
PGExplainer uses the following cross-entropy loss to train the explainer on multiple samples. This explainer acts as a global explainer by training a shared explanation model across multiple samples. Specifically, it learns an MLP that predicts edge importance scores. For binary classification, the objective is formulated as:
where is the original input graph for sample , is the sampled explanatory subgraph for sample in the -th sampling step, and represents the parameters of the explainer model. The architecture of PGExplainer is depicted in Figure 4.
@@
3.4 RankFusion Explainer: An Edge Ranking Aggregation Strategy
To enhance the quality of explanations across different sparsity levels, we propose a novel GNN explanation method called RankFusion explainer. This approach aggregates the edge rankings from two state-of-the-art explainers to generate more stable and accurate explanations. The process begins with constructing a bank of candidate explainers, including GNNExplainer, PGExplainer, CaptumExplainer, and other techniques. Each explainer in this bank is individually assessed by computing its accuracy across a range of sparsity levels. This is done by retaining the top-ranked edges, reconstructing the graph, and measuring the GNN’s prediction accuracy. Based on the overall performance, top two explainers are selected for fusion.
After the selection stage, the RankFusion algorithm transforms the edge importance scores from both explainers into descending rank orders. For each edge, the algorithm calculates the absolute difference between their assigned ranks. If this difference is smaller than a predefined threshold (expressed as a percentage of the total number of edges), the fused rank is conservatively assigned as the maximum of the two ranks. If the rank difference exceeds the threshold, the rank of that edge is taken from the explainer with higher accuracy at the corresponding explanation percentage, ensuring that disagreements are resolved in favor of the more reliable explainer.
Once all edges have been assigned fused rankings, the top-ranked edges based on the new weights are selected to reconstruct a sparsified version of the input graph. This reconstructed subgraph is then fed into the trained GNN model, and its prediction accuracy is recorded as the evaluation metric for the RankFusion explainer. This fusion-based strategy balances consensus and performance priority, leveraging the strengths of multiple explainers to improve robustness and fidelity in graph-based explanations across varying sparsity constraints. The detailed algorithmic steps of the RankFusion explainer are presented in Algorithm 2.
4 Explainer Subgraph Evaluation and Construction
@@
where is the important subgraph, is the original input, and is the trained GNN model and is the ground truth label. measures the prediction difference between the original input and its unimportant part, capturing the contribution of the removed edges to the model’s decision. Conversely, compares the prediction between the original graph and the important subgraph, reflecting the predictive capacity of the extracted subgraph. A higher value indicates that removing the identified important subgraph significantly alters the model’s prediction, confirming that this subgraph is crucial for the decision made. Therefore, higher scores correspond to more faithful explanations. In contrast, a lower value means that the important subgraph alone can nearly reproduce the model’s prediction on the original graph, suggesting it contains most of the essential information. Thus, lower values reflect higher-quality explanations.
@@
Each perturbed graph is retained for consistency evaluation if it meets the following two conditions. First, the model prediction must be unchanged:
Second, the graph-level embeddings must remain similar:
where denotes the embedding function of the model, and is a threshold for the similarity metric.
@@ 4.0.1 Top-Edge Selection (TES)
4.0.2 Greedy Edge-wise Composition (GEC)
@@
Algorithm 3 presents the GEC algorithm. By prioritizing both edge weight and structural connectivity, GEC ensures that the selected subgraph retains the most informative and influential components of the original graph. This results in more faithful and robust explanations, ultimately supporting a more reliable interpretation of the model’s decision-making process.
5 Results and Analysis
For our experiments, we selected malicious samples from BODMAS [30] and PMML [31], along with benign samples from DikeDataset [32]. Details of the datasets used in this study are summarized in Table 1. To recover CFGs dynamically, we employ the Angr library [33], a Python-based binary analysis tool. Angr constructs graphs by integrating both symbolic execution and constraint solving, enabling comprehensive and precise CFG recovery. The dynamically generated CFGs used in our experiments are available for public access111Available at http://cicresearch.ca/CICDataset/CIC-DGG-2025/Dataset/. Moreover, the implementation used in this study has been made publicly available222Available at https://github.com/GriffinUNB/On-the-Consistency-of-GNN-Explanations-for-Malware-Detection.
| Datasets | #Samples | Avg. Nodes | Avg. Edges | Class |
| BODMAS | 1117 | 57587.82 | 60252.08 | Malware |
| DikeDataset | 510 | 6011.56 | 10054.94 | Benign |
| PMML | 1029 | 11303.51 | 19073.31 | Malware |
We employed a symmetrical autoencoder to reduce the original 439-dimensional node feature vectors to a 64-dimensional latent space. The encoder comprises three fully connected layers with dimensions 439 256 128 64, each followed by a ReLU activation function. The decoder mirrors this structure with layers 64 128 256 439, also using ReLU activations. The autoencoder was trained for 5000 epochs using the Adam optimizer with a learning rate of 0.0001, minimizing the mean squared error between the input and reconstructed features. Training was stopped at 5000 epochs because the validation MSE plateaued below a threshold of for the final 1000 epochs, indicating convergence. The 64-dimensional embeddings produced by the encoder were subsequently used as the input node features for downstream graph learning tasks.
For graph classification, we experimented with several GNN architectures, including GCN, GAT, GraphSAGE, and GIN. All models share the same architecture: three graph convolution layers with 64 hidden units each, followed by ReLU activations. A global mean pooling layer aggregates node-level embeddings into a graph-level representation, which is then passed through a dropout layer (with a dropout rate of 0.2) and a final fully connected linear layer to produce class scores for binary classification. Each model was trained using the Adam optimizer with a learning rate of 0.0001 and weight decay of 0.0005, using the cross-entropy loss over 50 epochs. The dataset was randomly split into 50% for training and 50% for testing. Table 2 summarizes the class-wise evaluation metrics (precision, recall, and F1-score) for both benign and malicious classes, along with the overall accuracy for each model. Among the tested architectures, GCN outperformed the others, achieving the highest accuracy of 96.02%, along with a benign F1-score of 0.8927 and a malicious F1-score of 0.9756. The ROC curves for each model are presented in Figure 5, providing additional insight into their detection capabilities. Based on its superior overall performance, GCN was selected as the backbone model for the downstream explanation task.
| Model | Accuracy | Benign Class | Malicious Class | ||||
| Precision | Recall | F1-score | Precision | Recall | F1-score | ||
| GCN [10] | 0.9602 | 0.9498 | 0.8421 | 0.8927 | 0.9624 | 0.9891 | 0.9756 |
| GIN [11] | 0.7056 | 0.4000 | 0.9960 | 0.5708 | 0.9984 | 0.6347 | 0.7760 |
| GraphSAGE [12] | 0.9515 | 0.8577 | 0.9028 | 0.8797 | 0.9759 | 0.9634 | 0.9696 |
| GAT [13] | 0.9523 | 0.8912 | 0.8623 | 0.8765 | 0.9666 | 0.9743 | 0.9704 |
@@ 5 Results and Analysis
We assess the performance of the explainers using the , , and Consistency metrics across different sparsity levels. Figure 7 presents the and results under the GEC approach. In each subplot, the x-axis represents the sparsity level, defined as the percentage of edges retained from the original graph, and the y-axis indicates the fidelity score. The solid lines show the fidelity values computed from the original, unperturbed graphs for each explainer and sparsity level.
@@
To assess consistency in practice, we generated 20 perturbed versions of each input graph: 10 with random node removal and 10 with random edge removal. In our implementation, the perturbation ratio was set to . Each perturbed graph was retained for analysis only if it preserved the model’s original prediction and the cosine similarity between the graph-level embeddings of the original and perturbed graphs was below a selected threshold. We performed a sweep over similarity thresholds ranging from 0.01 to 0.03 in increments of 0.0025 to identify valid perturbations. For each explainer, we visualized the variability of and scores across all valid perturbations. The solid lines represent fidelity values computed on the original (unperturbed) graphs, while the shaded regions reflect the range between the minimum and maximum fidelity scores observed across the valid perturbations.
@@
The results show that values are generally low and stable for explainers, when subgraphs are extracted using GEC. This is particularly evident for GNNExplainer, whose performance stabilizes significantly under this approach. According to the definition of , lower values are desirable, and this trend is clearly observed for the best-performing explainers (IG, GBP, and SAL), which consistently yield lower scores.
@@
In contrast, values demonstrate a consistent pattern across sparsity levels. The top explainers (IG, GBP, and SAL) achieve higher scores, indicating a substantial change in model prediction when the important subgraph is removed. This suggests that the identified important subgraph indeed contains the critical predictive structure. Furthermore, the consistency bands in the plot are generally narrow, indicating that explanations remain stable across perturbations. Although bands are somewhat wider, the results still indicate reasonable stability, particularly for the top explainers. These findings highlight the advantage of using GEC in conjunction with effective explainers for generating reliable and robust explanations.
@@
To assess the effectiveness of the proposed RankFusion explainer, we investigate whether selecting and combining the outputs of the top-performing explainers can lead to improved subgraph quality and model performance. Based on the previous analysis across accuracy, fidelity, and consistency metrics, the two most effective explainers identified were Integrated Gradients and Guided Backpropagation. RankFusion is applied by aggregating the edge importance rankings produced by these two explainers using the methodology described earlier. To provide a more comprehensive evaluation, we also compare RankFusion against two additional aggregation baselines: mean aggregation and rank voting. For mean aggregation, the normalized edge importance scores of IG and GBP are averaged. For rank voting, three explainers (IG, GBP, and Saliency) are considered, and edges are selected based on a majority vote (i.e., 2 out of 3 agreement) among their top-k ranked edges.
@@
Figure 8 compares the model’s accuracy when using subgraphs generated by IG, GBP, and their aggregated result. All subgraphs in this evaluation are extracted using the GEC method. The x-axis represents the sparsity level, and the y-axis shows the accuracy obtained by feeding the resulting subgraph into the trained GCN. Across all sparsity levels, RankFusion consistently outperforms both the individual explainers and the alternative aggregation methods. While the improvement over the best individual explainer (IG) is modest, the results clearly support the hypothesis that selecting and combining the outputs of strong explainers leads to more informative subgraphs.
@@
6 Conclusion
This paper presents a comprehensive framework for malware detection using dynamically generated CFGs. Node-level assembly instructions are embedded into a continuous vector space through an embedding approach that combines rule-based encoding with autoencoder-based learning. These representations are then used as input to a GNN-based classifier for malware detection. To enhance the interpretability of the model, five explainers are employed: GNNExplainer, PGExplainer, and CaptumExplainer with three attribution methods, namely Integrated Gradients, Guided Backpropagation, and Saliency. In addition, we propose RankFusion, an aggregated explainer that selects and integrates the outputs of the top-performing explainers to improve the quality of generated explanations. We also introduce GEC as a connectivity-aware subgraph extraction strategy that mitigates the shortcomings of traditional top-edge selection. Experimental results demonstrate the strong detection performance of the framework, with an accuracy exceeding 94% and an F1 score above 96%. Further analysis based on fidelity, accuracy, and consistency metrics shows that the combination of RankFusion and GEC produces more stable, interpretable, and accurate explanations, reinforcing the value of incorporating both explainer aggregation and structurally coherent subgraph extraction in GNN-based malware detection. Future work will explore extending the framework beyond binary classification to identify specific malware families, enabling finer-grained threat analysis. Additionally, incorporating temporal features alongside CFGs could enrich the model’s understanding of malware dynamics. Another promising direction involves investigating the robustness of both the classifier and the explainers against adversarial perturbations, which is critical for maintaining reliability in security-sensitive applications.